OpenStudy (anonymous):

During a research experiment, it was found that the number of bacteria in a culture grew at a rate proportional to its size. At 7:00AM there were 5000 bacteria present in the culture. At noon the number of bacteria grew to 5,900. How many bacteria will there be at midnight?

OpenStudy (anonymous):

it seems as if every 5 hours 900 bacteria grow. so 900/5= 180 that means every hour 180 more bacteria will grow. so from noon to midnite how many hours pass?

OpenStudy (anonymous):

that how I did it in the test.900/5=180*12=2160+5900=8060.But when I got the test back the answer was 8777 and I dont know why?

OpenStudy (anonymous):

I'm not sure maybe you should ask the teacher.